OpenStudy (anonymous):

Solve the system of equations. 8x + 5y = 4 3x + 2y = 3 a.(3, -4) b.(-7, 12) c.(1, 0) d.(5, 6) don't understand system of equations, pleasee help!

OpenStudy (anonymous):

b

mathslover (mathslover):

ok : let \(\huge{8x+5y=4}\) ----------- 1) and \(\huge{3x+2y=3}\) --------------2) now solve for x in the 1st equation : 1) \[\huge{8x+5y=4}\] \[\huge{8x=4-5y}\] \[\huge{x=\frac{4-5y}{8}}\] now put this value in the 2nd equation for x can u solve that ?

OpenStudy (anonymous):

If you do not know how to deal with, simply substitute the answers to the equation)

OpenStudy (anonymous):

yes let me try quick

mathslover (mathslover):

no problem try ur best

OpenStudy (anonymous):

i got (-7,12) am I correct

mathslover (mathslover):

great ....